OpenStudy (anonymous):

I have like 7 problems I need help with but I would have to email them because theyre are graphs attatched...does anyone feel like being a HUGE help>>??

OpenStudy (anonymous):

You can attach files here. Just do 'em one by one.

OpenStudy (anonymous):

which of the following is the equation of the horizontal asymptote of the graph of f(x) = (1/3)^(x+2) -3 x = -2 x=0 y=1/3 y=-3

OpenStudy (anonymous):

y=-3

OpenStudy (anonymous):

To solve this, firstly, it asks for a horizontal asymptote, so the first two are out as they would give vertical ones. Next you need to set y=-3 in the equation, and so 0=(1/3)^(x+2), as you have to take logs to solve for x, and ln(0) is not defined, y=-3 is not defined Hope this helps